OpenStudy (anonymous):

Use the first derivative to determine where the function f(x)=5x^4+1280x is increasing and decreasing

OpenStudy (anonymous):

f(x) = 5x^4 + 1280x f'(x) = 20x^3 + 1280 f'(x) = 0 = 20x^3 + 1280 -1280 = 20x^3 -64 = x^3 x = -4

OpenStudy (anonymous):

This value is a critical point, so you just need to test on either side of this x-value to determine increasing (positive value) and decreasing (negative value) parts of the interval.